OpenStudy (anonymous):

complete the table value for the operation 6x+3y=36 x y 0 0 3 Can someone help me with this please

OpenStudy (anonymous):

what do you get when you plug 0 in for x in the equation

OpenStudy (anonymous):

6*0 i think

OpenStudy (anonymous):

like 6x+3y=36 so 6*0+3y=36 and I am lost after that

OpenStudy (anonymous):

right so that simplifies to 0 + 3y= 36 0 +3y is the same thing as 3y so 3y = 36 divide by 3 to get y by itself so y = 36/3 y = 12

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes so would i do the same if the x is 3

OpenStudy (anonymous):

yup! and for teh middle option plug in 0 for y and solve for x:)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

ok so I tried this an came up with 6 divided by 36=6 for x

OpenStudy (anonymous):

so u got 6 for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good job:))

OpenStudy (anonymous):

what did u get for teh last one